If anyone could please answer this question i would thank you graciously. g(x)= 5sqrt(x+1)-7
Answer by stanbon(48535)

g(x)= 5sqrt(x+1)-7
-------------------
Determine the inverse.
y = 5sqrt(x+1)-7
Interchange x and y as follows:
x = 5sqrt(y+1)-7
Solve for y as follows:
5sqrt(y+1) = x+7
sqrt(y+1) = (x+7)/5
y+1 = [(x+7)/5]^2
y = [(x+7)/5]^2 -1
That is the inverse.
=======================
Cheers,
Stan H.